# Confluent Vandermonde matrix and related topics

### v2

@@ Abstract
In this note, we explore the connections between the confluent Vandermonde matrix and several mathematical topics, including interpolating polynomials, Hasse derivatives, LU factorization, companion matrices and their Jordan forms, and the Chinese remainder theorem. Using a unified approach based on polynomial evaluations and derivative computations at selected points, we provide accessible proofs that not only clarify key results but also offer insights for both experienced researchers and those new to the subject.
Keywords: Confluent Vandermonde matrix, interpolating polynomials, Hasse derivatives, LU factorization, companion matrix, Jordan form, Chinese remainder theorem.
@@ 1 Introduction
The Vandermonde matrix has been used in polynomial interpolation, coding theory, signal processing, differential equations, and recursive relations; see [20, 14, 5, 2, 16] and the references therein. It was also used to give alternative proofs for some classical results. For example, it was used to prove a theorem of Abel [11], whose classical proof is based on the residue theorem. There are many elegant results on the Vandermonde matrix, and one of them is the following formula for the determinant of $\det(V_{\Lambda})$ which appeared in many textbooks
$\det(V_{\Lambda})=\prod_{i<j}(\lambda_{j}-\lambda_{i}).$
Rushanan[20] gave the LU factorization of $V_{\Lambda}$. Many formulas for $V_{\Lambda}^{-1}$ were studied and rediscovered by researchers in pure and applied subjects; see, e.g., [17, 15]. In particular, Parker[19] pointed out that the columns of $V^{-1}$ are exactly the coefficients of the Lagrange polynomials. The generalization of $V_{\Lambda}$ whose columns are assigned with nonconsecutive powers was considered and revealed a strong connection to symmetric functions through Schur functions; see, e.g., [18, 12, 10, 7, 3, 6].
Brand[1] pointed out an interesting relation that $V_{\Lambda}C_{\Lambda}V_{\Lambda}^{-1}$ is a diagonal matrix, where $C$ is the companion matrix of $(x-\lambda_{1})\cdots(x-\lambda_{d})$. In the same paper, Brand mentioned that if $\lambda_{1},\dots,\lambda_{d}$ are not distinct, one can extend the definition of the Vandermonde matrix to define the confluent Vandermonde matrix $V_{\Lambda}$ using the higher derivatives of the polynomial $(x-\lambda_{1})\cdots(x-\lambda_{d})$ at the points $\lambda_{j}$, which appear more than once; see the definition in Section 2. Then $V_{\Lambda}C_{\Lambda}V_{\Lambda}^{-1}$ will be the Jordan canonical form of $C_{\Lambda}$.
@@ 2 Evaluations of a polynomial at specific points
Therefore, $V_{\Lambda}$ can be viewed as a transformation of a polynomial of degree at most $d-1$ to its evaluations at the points in $\Lambda$. In fact, $V_{\Lambda}$ is a linear transformation from the linear space of polynomials with degrees bounded by $d-1$ to ${\mathbb{R}}^{d}$. Since $\det(V_{\Lambda})\neq 0$, $V_{\Lambda}$ is invertible and any assignment of $p(\lambda_{1}),\ldots,p(\lambda_{d})$ defines a unique polynomial $p(x)$ of degree bounded by $d-1$.
@@
$\begin{array}[]{lcrrrr}\frac{1}{0!}\cdot p(x)&=&c_{0}+&c_{1}x+&c_{2}x^{2}+&\cdots+c_{d-1}x^{d-1},\\ \frac{1}{1!}\cdot\frac{d}{dx}p(x)&=&&c_{1}+&c_{2}2x+&\cdots+c_{d-1}(d-1)x^{d-2},\\ \frac{1}{2!}\cdot\frac{d^{2}}{dx^{2}}p(x)&=&&&c_{2}+&\cdots+c_{d-1}\left(\frac{(d-1)(d-2)}{2}\right)x^{d-3},\\ &\vdots&&&&\end{array},$
@@
$V_{\Lambda}^{(d)}=\begin{bmatrix}-&V_{\lambda_{1}^{(m_{1})}}^{(d)}&-\\ &\vdots&\\ -&V_{\lambda_{q}^{(m_{q})}}^{(d)}&-\\ \end{bmatrix}.$
The square matrix $V_{\Lambda}=V_{\Lambda}^{(d)}$ is called the confluent Vandermonde matrix (a.k.a. the generalized Vandermonde matrix) of $\Lambda$.
@@
Let $[p(x)]=[c_{0},\dots,c_{d-1}]^{\top}$ be the vector of coefficients of $p(x)=c_{0}+c_{1}x+\cdots+c_{d-1}x^{d-1}$ and $\operatorname{eval}_{\Lambda}(p(x))$ the the evaluations of $p(x)$ as shown in the right-hand side of Equation 5. Hence, we have $V_{\Lambda}[p(x)]=\operatorname{eval}_{\Lambda}(p(x))$ extending the classical result Equation 3.
@@ 3 Hasse derivative and confluent Vandermonde matrices over arbitrary fields
Note that the binomial coefficient $\binom{n}{k}$ is the coefficient of $x^{k}$ in the expansion of $(1+x)^{n}$, which is well defined in any field and is the same as $\frac{n!}{k!(n-k)!}$ when $\mathbb{F}$ is of characteristic 00. For this sake, an alternative definition of the Hasse derivative is $\mathcal{D}^{(r)}p(x)=[y^{r}]p(x+y)$, which is the coefficient of $y^{r}$ in $p(x+y)$ when treated as a polynomial of $y$. This can be verified by direct computation as follows.
@@ 4 LU factorization
Let $U^{-1}=\Big{[}[p_{0}(x)]\ \cdots\ [p_{d-1}(x)]\Big{]}$ and $L=\Big{[}\operatorname{eval}_{\Lambda}(p_{0}(x))\ \cdots\ \operatorname{eval}_{\Lambda}(p_{d-1}(x))\Big{]}$. Note that $U^{-1}$ is an upper triangular matrix with $1$’s on the diagonal. Thus, we can write it as the inverse of some matrix $U$, where $U=(U^{-1})^{-1}$ is an upper triangular matrix. Moreover, by Corollary 3.2, $L$ is a lower triangular matrix with the diagonal entries $1$ for $m_{1}$ times, $(\lambda_{2}-\lambda_{1})^{m_{1}}$ for $m_{2}$ times, $(\lambda_{3}-\lambda_{1})^{m_{1}}(\lambda_{3}-\lambda_{2})^{m_{2}}$ for $m_{3}$ times, and so on. Therefore, we have the following.
@@
Let $\Lambda=\{1^{(1)},3^{(2)},6^{(2)}\}$. Then
@@
Since $\det(U^{-1})=1$, it follows that $\det(V_{\Lambda})=\det(L)=(3-1)^{2}(6-1)^{2}(6-3)^{4}$.
5 The Chinese Remainder Theorem
Let $\Lambda=\{\lambda_{1}^{(m_{1})},\ldots,\lambda_{q}^{(m_{q})}\}$ with $d=m_{1}+\cdots+m_{q}$. Define
@@ 5 The Chinese Remainder Theorem
for $j=1,\ldots,q$ and $m=0,\ldots,m_{j}-1$. Note that there are $d$ polynomials. When we write $\hat{h}_{1,0}(x)$, $\ldots$, $\hat{h}_{q,m_{q}-1}(x)$, we mean $\hat{h}_{1,0}(x)$, $\ldots$, $\hat{h}_{1,m_{1}-1}(x)$, $\ldots$, $\hat{h}_{q,0}(x)$, $\ldots$, $\hat{h}_{q,m_{q}-1}(x)$ in order. The Chinese Remainder Theorem for polynomial rings asserts that every polynomial of degree less than $d$ can be uniquely written as a linear combination of ^$\hat{\mathcal{H}}_{\Lambda}=\{\hat{h}_{1,0}(x),\ldots,\hat{h}_{q,m_{q}-1}(x)\}$; equivalently, ^$\hat{\mathcal{H}}_{\Lambda}$ is a basis of the space of all polynomial of degree less than $d$. For example, using Example 2.2, we have the following.
@@
Let $\Lambda=\{\lambda_{1}^{(1)},\lambda_{2}^{(2)},\lambda_{3}^{(2)}\}$. Then
$\hat{h}_{1,0}(x)=(x-\lambda_{2})^{2}(x-\lambda_{3})^{2},\hskip 1.5pt\in\ ~{}~{}~{}$
$\hat{h}_{2,0}(x)=(x-\lambda_{1})(x-\lambda_{3})^{2},\ \hat{h}_{2,1}(x)=(x-\lambda_{2})(x-\lambda_{1})(x-\lambda_{3})^{2},$
$\hat{h}_{3,0}(x)=(x-\lambda_{1})(x-\lambda_{2})^{2},\ \hat{h}_{3,1}(x)=(x-\lambda_{3})(x-\lambda_{1})(x-\lambda_{2})^{2}$
form a basis for the linear space of polynomial of degree less than 5.
The confluent Vandermonde matrix provides a simple way to prove this version of the Chinese Remainder Theorem as follows. Let
Then the Chinese Remainder Theorem is equivalent to that $\hat{H}$ is invertibile. To see this, we observe that
$\mathcal{D}^{(r)}\hat{h}_{j,m}(\lambda_{i})=\begin{cases}0&\text{if }i\neq j\text{ and }r<m_{i},\\ 0&\text{if }i=j\text{ and }r<m,\\ \star&\text{if }i=j\text{ and }r=m,\\ ?&\text{if }i=j\text{ and }r>m.\\ \end{cases}$
Here $\star$ stands for a nonzero number and ???? stands for any value in the field. To be more precise, by Corollary 3.2 we have
if $\lambda_{i}-\lambda_{j}\neq 0$ whenever $i\neq j$. Consequently, $V_{\Lambda}\hat{H}$ is a lower triangular matrix with diagonal entries all nonzero. Therefore, both $V_{\Lambda}$ and $\hat{H}$ are invertible, proving the Chinese Remainder Theorem.
Note that this argument also proves the invertibility condition of $V_{\Lambda}$ without using the formula in Theorem 4.2.
Again, using Example 2.2 with $(\lambda_{1},\lambda_{2},\lambda_{3})=(1,3,6)$, we have the following.
Example 5.2.
Let $\Lambda=\{1^{(1)},3^{(2)},6^{(2)}\}$. Then
@@
We may compute that
@@
$V_{\Lambda}=\begin{bmatrix}1&1&1&1&1\\ 1&3&9&27&81\\ 0&1&6&27&108\\ 1&6&36&216&1296\\ 0&1&12&108&864\end{bmatrix}\text{ and }\hat{H}_{\Lambda}=\begin{bmatrix}324&-36&108&-9&54\\ -324&48&-180&15&-99\\ 117&-13&87&-7&57\\ -18&1&-16&1&-13\\ 1&0&1&0&1\end{bmatrix}.$
Also,
$V_{\Lambda}\hat{H}_{\Lambda}=\begin{bmatrix}\hat{h}_{1,0}(1)&\cdots&\hat{h}_{3,1}(1)\\ \hat{h}_{1,0}(3)&\cdots&\hat{h}_{3,1}(3)\\ \mathcal{D}^{(1)}\hat{h}_{1,0}(3)&\cdots&\mathcal{D}^{(1)}\hat{h}_{3,1}(3)\\ \hat{h}_{1,0}(6)&\cdots&\hat{h}_{3,1}(6)\\ \mathcal{D}^{(1)}\hat{h}_{1,0}(6)&\cdots&\mathcal{D}^{(1)}\hat{h}_{3,1}(6)\\ \end{bmatrix}=\begin{bmatrix}100&0&0&0&0\\ 0&18&0&0&0\\ 0&-3&18&0&0\\ 0&0&0&45&0\\ 0&0&0&39&45\end{bmatrix}.$
@@
Let $\Lambda=\{\lambda_{1}^{(m_{1})},\dots,\lambda_{q}^{(m_{q})}\}$ with $m_{1}+\cdots+m_{q}=d$. Then $T_{\Lambda}=V_{\Lambda}\hat{H}_{\Lambda}$ is lower triangular matrix and is a block-diagonal matrix with $q$ blocks. For any given polynomial $p(x)$ of degree less than $d$, $\hat{H}^{-1}_{\Lambda}[p(x)]=T^{-1}_{\Lambda}V_{\Lambda}[p(x)]$ records the coefficients of $p(x)$ as a linear combination of ^$\hat{\mathcal{H}}_{\Lambda}$.
6 Hermite polynomials

### v3

@@ Abstract
In this note, we explore the connections between the confluent Vandermonde matrix over an arbitrary field and several mathematical topics, including interpolation polynomials, Hasse derivatives, LU factorization, companion matrices and their Jordan forms, and the partial fraction decomposition. Using a unified approach based on polynomial evaluations and derivative computations at selected points, we provide accessible proofs that not only clarify key results but also offer insights for both experienced researchers and those new to the subject.
Keywords: Confluent Vandermonde matrix, interpolation polynomials, Hasse derivatives, LU factorization, companion matrix, Jordan form, partial fraction decomposition.
@@ 1 Introduction
The Vandermonde matrix has been used in polynomial interpolation, coding theory, signal processing, differential equations, and recursive relations; see [25, 16, 7, 4, 18] and the references therein. It was also used to give alternative proofs for some classical results. For example, it was used to prove a theorem of Abel [13], whose classical proof is based on the residue theorem. There are many elegant results on the Vandermonde matrix, and one of them is the following formula for the determinant of $V_{\Lambda}$ which appeared in many textbooks
$\det(V_{\Lambda})=\prod_{i<j}(\lambda_{j}-\lambda_{i}).$ (2)
Rushanan[25] gave the LU factorization of $V_{\Lambda}$. Many formulas for $V_{\Lambda}^{-1}$ were studied and rediscovered by researchers in pure and applied subjects; see, e.g., [19, 17]. In particular, Parker[22] pointed out that the columns of $V_{\Lambda}^{-1}$ are exactly the coefficients of the Lagrange interpolation polynomials. The generalization of $V_{\Lambda}$ whose columns are assigned with nonconsecutive powers was considered and revealed a strong connection to symmetric functions through Schur functions; see, e.g., [21, 14, 12, 9, 5, 8, 1].
Using an example, Aitken and Turnbull [2, Chapter VI] demonstrated an interesting fact that $V_{\Lambda}C_{\Lambda}V_{\Lambda}^{-1}$ is a diagonal matrix, where $C_{\Lambda}$ is the companion matrix of $(x-\lambda_{1})\cdots(x-\lambda_{d})$. In the same chapter, the authors mentioned that if $\lambda_{1},\dots,\lambda_{d}$ are not distinct, one can extend the definition of the Vandermonde matrix to define the confluent Vandermonde matrix $V_{\Lambda}$ using the higher derivatives of the polynomial $(x-\lambda_{1})\cdots(x-\lambda_{d})$ at the points $\lambda_{j}$, which appear more than once; see the definition in Section 2. Then $V_{\Lambda}C_{\Lambda}V_{\Lambda}^{-1}$ will be the Jordan canonical form of $C_{\Lambda}$. The same points were later emphasized in Brand [3].
@@ 2 Evaluations of a polynomial at specific points
Therefore, $V_{\Lambda}$ can be viewed as a transformation from a polynomial of degree at most d1d-1 to its evaluations at the points in $\Lambda$. In fact, $V_{\Lambda}$ is a linear transformation from the linear space of polynomials with degrees bounded by d1d-1 to ${\mathbb{R}}^{d}$. Since $\det(V_{\Lambda})\neq 0$, $V_{\Lambda}$ is invertible and any assignment of $p(\lambda_{1}),\ldots,p(\lambda_{d})$ defines a unique polynomial $p(x)$ of degree bounded by d1d-1.
@@
[…]
@@
Kalman [16] gave a detailed discussion on this approach, while the confluent Vandermonde matrix to be defined below can be traced back to Aitken and Turnbull [2] and Aitken [1].
@@
$V_{\Lambda}=\begin{bmatrix}-&V_{\lambda_{1}^{(m_{1})}}^{(d)}&-\\ &\vdots&\\ -&V_{\lambda_{q}^{(m_{q})}}^{(d)}&-\\ \end{bmatrix}$.
The square matrix $V_{\Lambda}$ is called the confluent Vandermonde matrix (a.k.a. the generalized Vandermonde matrix) of $\Lambda$.
@@
Let $p(x)]=[c_{0},\dots,c_{d-1}]^{\top}$ be the vector of coefficients of $p(x)=c_{0}+c_{1}x+\cdots+c_{d-1}x^{d-1}$ and $\operatorname{eval}_{\Lambda}(p(x))$ the the evaluations of $p(x)$ as shown in the right-hand side of Equation 6. Hence, we have $V_{\Lambda}[p(x)]=\operatorname{eval}_{\Lambda}(p(x))$ extending the classical result Equation 4.
@@ 3 Hasse derivative and confluent Vandermonde matrices over arbitrary fields
Note that the binomial coefficient $\binom{n}{k}$ is the coefficient of $x^{k}$ in the expansion of $(1+x)^{n}$, which is well defined in any field and is the same as $\frac{n!}{k!(n-k)!}$ when $\mathbb{F}$ is of characteristic 0. For this sake, an alternative definition of the Hasse derivative is $\mathcal{D}^{(r)}p(x)=[y^{r}]p(x+y)$, which is the coefficient of $y^{r}$ in $p(x+y)$ when treated as a polynomial of $y$. This can be verified by direct computation as follows.
@@ 4 LU factorization
Let $U^{-1}=\Big{[}[p_{0}(x)]\ \cdots\ [p_{d-1}(x)]\Big{]}$ and $L=\Big{[}\operatorname{eval}_{\Lambda}(p_{0}(x))\ \cdots\ \operatorname{eval}_{\Lambda}(p_{d-1}(x))\Big{]}$. Note that $U^{-1}$ is an upper triangular matrix with $1$’s on the diagonal. Thus, we can write it as the inverse of some matrix $U$, where $U=(U^{-1})^{-1}$ is an upper triangular matrix. Moreover, by Corollary 3.2, $L$ is a lower triangular matrix with the diagonal entries $1$ for $m_{1}$ times, $(\lambda_{2}-\lambda_{1})^{m_{1}}$ for $m_{2}$ times, $(\lambda_{3}-\lambda_{1})^{m_{1}}(\lambda_{3}-\lambda_{2})^{m_{2}}$ for $m_{3}$ times, and so on.
This approach was used by Rushanan[25] to develop the LU factorization of $V_{\Lambda}$ when all elements in $\Lambda$ are distinct and $V_{\Lambda}$ is the classical Vandermonde matrix. As we have seen, the same approach leads to the similar consequences for the confluent Vandermonde matrix. Therefore, we have the following.
@@
This leads to a generalization of Equation 2. The next theorem was mentioned in [2] and proved in [1]. The proof in [1] requires arithmetic on homogeneous symmetric functions and taking limits when $\mathbb{F}=\mathbb{R}$, while using the LU factorization allows us to observe the formula directly for $V_{A}$ over arbitrary $\mathbb{F}$.
@@
Let $\Lambda=\{1^{(1)},3^{(2)},6^{(2)}\}$ and $\mathbb{F}$ a field where $1,3,6$ are distinct, or equivalently, $\operatorname{char}(\mathbb{F})\neq 2,3,5$. Then
@@
Since $\det(U^{-1})=1$, it follows that $\det(V_{\Lambda})=\det(L)=(3-1)^{2}(6-1)^{2}(6-3)^{4}$. Since $1,3,6$ are distinct in $\mathbb{F}$, we have $\det(L)\neq 0$.
Remark 4.4.
The Chinese remainder theorem states that given the residues of a polynomial modulo $p_{i}(x)$ for some coprime family of polynomials $\{p_{1}(x),\ldots,p_{q}(x)\}$, the polynomial always exists and is uniquely determined, modulo $\prod_{i=1}^{q}p_{i}(x)$; see, e.g., [26]. When $p_{i}(x)=(x-\lambda_{i})^{m_{i}}$ for some multiset $\Lambda=\{\lambda_{1}^{(m_{1})},\ldots,\lambda_{q}^{(m_{q})}\}$, the Chinese remainder theorem is equivalent to the invertibility of $V_{\Lambda}$.
5 Partial fraction decomposition
Partial fraction decomposition is an important tool for many areas, e.g., calculus, complex analysis, and more. However, when the denominator has repeated roots, the computation of the partial fraction decomposition is not so straightforward; see, e.g., [20].
For example, let $\Lambda=\{1^{(1)},3^{(2)},6^{(2)}\}$. Given a polynomial $p(x)$ of degree at most $4$, how to find the coefficients $a,b,c,d,e$ such that
$\frac{p(x)}{(x-1)(x-3)^{2}(x-6)^{2}}=\frac{a}{(x-1)}+\frac{b+c(x-3)}{(x-3)^{2}}+\frac{d+e(x-6)}{(x-6)^{2}}$ (8)
holds?
For the real and complex field, when a root $\lambda$ is simple, one may multiply on both sides by $(x-\lambda)$ and take the limit $x\rightarrow\lambda$ to find the corresponding coefficients. For example,
$a=p(1)\lim_{x\rightarrow 1}\frac{(x-1)}{(x-1)(x-3)^{2}(x-6)^{2}}.$
However, the same idea does not extend to the case for repeated roots and polynomials over arbitrary fields. Here we demonstrate an efficient approach using the evaluation $\operatorname{eval}(p(x))$ and the Vandermonde matrix $V_{\Lambda}$, which is similar to the methods in [20].
We first formulate the problem. Let $\Lambda=\{\lambda_{1}^{(m_{1})},\ldots,\lambda_{q}^{(m_{q})}\}$ with $d=m_{1}+\cdots+m_{q}$. Define
@@ 5 Partial fraction decomposition
for $j=1,\ldots,q$ and $m=0,\ldots,m_{j}-1$. Note that there are $d$ polynomials. When we write $\hat{h}_{1,0}(x)$, $\ldots$, $\hat{h}_{q,m_{q}-1}(x)$, we mean $\hat{h}_{1,0}(x)$, $\ldots$, $\hat{h}_{1,m_{1}-1}(x)$, $\ldots$, $\hat{h}_{q,0}(x)$, $\ldots$, $\hat{h}_{q,m_{q}-1}(x)$ in order. By writing Equation 8 in a common denominator, we observe that finding the partial fraction decomposition is the same as writing $p(x)$ as a linear combination of $\hat{\mathcal{H}}_{\Lambda}=\{\hat{h}_{1,0}(x),\ldots,\hat{h}_{q,m_{q}-1}(x)\}$. Let
Then coefficients of the partial fraction decomposition can be found by $\hat{H}_{\Lambda}^{-1}[p(x)]$.
@@
Let $\Lambda=\{1^{(1)},3^{(2)},6^{(2)}\}$ and $\mathbb{F}$ a field where $1,3,6$ are distinct. Then
@@
By expanding all polynomials, we have
$\hat{H}_{\Lambda}=\begin{bmatrix}324&-36&108&-9&54\\ -324&48&-180&15&-99\\ 117&-13&87&-7&57\\ -18&1&-16&1&-13\\ 1&0&1&0&1\end{bmatrix}$.
Thus, solving the coefficients in Equation 8 is equivalent to solving
$\hat{H}_{\Lambda}\begin{bmatrix}a\\ b\\ c\\ d\\ e\end{bmatrix}=[p(x)].$
However, it is time consuming to find the inverse of $\hat{H}_{\Lambda}$.
The confluent Vandermonde matrix $V_{\Lambda}$ makes the computation easier. Suppose we want to solve $\hat{H}_{\Lambda}\mathbf{x}=[p(x)]$ for $\mathbf{x}$. By multiplying $V_{\Lambda}$ to the both sides, we have $(V_{\Lambda}\hat{H}_{\Lambda})\mathbf{x}=\operatorname{eval}(p(x))$. We will see that $T_{\Lambda}=V_{\Lambda}\hat{H}_{\Lambda}$ is invertible and has nice structure, so that the answer can be readily determined by $\mathbf{x}=T_{\Lambda}^{-1}\operatorname{eval}(p(x))$.
Since $V_{\Lambda}$ transforms a polynomial into its evaluations, we examine the structure of $V_{\Lambda}\hat{H}_{\Lambda}$ by taking the evaluations of $\hat{h}_{j,m}(x)$. We observe that
[…]
Here $\star$ stands for a nonzero number and ?? stands for any value in the field. To be more precise, by Corollary 3.2 we have
if $\lambda_{i}-\lambda_{j}\neq 0$ whenever $i\neq j$. Moreover, Corollary 3.2 also guarantees $\mathcal{D}^{(r)}\hat{h}_{j,m}(\lambda_{j})=\mathcal{D}^{(r-1)}\hat{h}_{j,m-1}(\lambda_{j})$ if $r\geq 1$ and $m\geq 1$. Remark 5.2 summarizes all these properties.
Remark 5.2.
For any multiset $\Lambda$, the matrix $T_{\Lambda}=V_{\Lambda}\hat{H}_{\Lambda}$ is a lower triangular matrix with nonzero diagonal entries, so $T_{\Lambda}$, $V_{\Lambda}$, and $\hat{H}_{\Lambda}$ are all invertible, which shows the existence and the uniqueness of the partial fraction decomposition. Moreover, $T_{\Lambda}$ is the direct sum of $q$ Toeplitz matrices, where $q$ is the number of distinct values in $\Lambda$.
Continuing Example 5.1, we solve the partial fraction decomposition through $T_{\Lambda}=V_{\Lambda}\hat{H}_{\Lambda}$.
Example 5.3.
We first compute $T=V_{\Lambda}\hat{H}_{\Lambda}$ through evaluations. By direct computation, $\mathcal{D}^{(0)}\hat{h}_{1,0}(1)=100$. Also, we may compute that
@@
$T=V_{\Lambda}\hat{H}_{\Lambda}=\begin{bmatrix}\hat{h}_{1,0}(1)&\cdots&\hat{h}_{3,1}(1)\\ \hat{h}_{1,0}(3)&\cdots&\hat{h}_{3,1}(3)\\ \mathcal{D}^{(1)}\hat{h}_{1,0}(3)&\cdots&\mathcal{D}^{(1)}\hat{h}_{3,1}(3)\\ \hat{h}_{1,0}(6)&\cdots&\hat{h}_{3,1}(6)\\ \mathcal{D}^{(1)}\hat{h}_{1,0}(6)&\cdots&\mathcal{D}^{(1)}\hat{h}_{3,1}(6)\\ \end{bmatrix}=\begin{bmatrix}100&0&0&0&0\\ 0&18&0&0&0\\ 0&-3&18&0&0\\ 0&0&0&45&0\\ 0&0&0&39&45\end{bmatrix}$.
By the structure of $T_{\Lambda}$, its inverse can be easily found as
$T^{-1}_{\Lambda}=\begin{bmatrix}\frac{1}{100}&0&0&0&0\\ 0&\frac{1}{18}&0&0&0\\ 0&\frac{1}{108}&\frac{1}{18}&0&0\\ 0&0&0&\frac{1}{45}&0\\ 0&0&0&-\frac{13}{675}&\frac{1}{45}\end{bmatrix}$.
Note that $\operatorname{char}(\mathbb{F})\neq 2,3,5$, the entries in $T_{\Lambda}^{-1}$ are well defined. With $\operatorname{eval}(p(x))$ computed, the coefficients of partial fraction decomposition of $p(x)$ are recorded in $T^{-1}_{\Lambda}\operatorname{eval}(p(x))$.
@@
Let $\Lambda=\{\lambda_{1}^{(m_{1})},\dots,\lambda_{q}^{(m_{q})}\}$ with $m_{1}+\cdots+m_{q}=d$. For any given polynomial $p(x)$ of degree at most d1d-1, $\hat{H}^{-1}_{\Lambda}[p(x)]=T^{-1}_{\Lambda}\operatorname{eval}(p(x))$ records the coefficients of $f(x)$ as a linear combination of $\hat{\mathcal{H}}_{\Lambda}$, which gives the partial fraction decomposition of $f(x)/\prod_{i=1}^{q}(x-\lambda_{i})^{m_{i}}$.
[…]
Figure 1: Relations between $[p(x)]$, $\operatorname{eval}(p(x))$, and the partial fraction decomposition.
In conclusion, Figure 1 illustrates the relations between $[p(x)]$, $\operatorname{eval}(p(x))$, and the partial fraction decomposition. With $\Lambda$ given, each of $V_{\Lambda}$, $\hat{H}_{\Lambda}$, and $T_{\Lambda}$ can be computed directly. By Remark 5.2, it is not too hard to find $T_{\Lambda}^{-1}$. Thus, Theorem 5.4 states that $\hat{H}_{\Lambda}^{-1}=T_{\Lambda}^{-1}V_{\Lambda}$ provides a more efficient way to compute the partial fraction decomposition.
6 Hermite interpolation polynomials
Given the evaluations of a polynomial at different points, the Lagrange interpolation polynomials provide a powerful tool in reconstructing the polynomial. What happens if the evaluations have derivatives involved?